# CONDA 2024 Shared Task on Evidence of Data Contamination: First Compilation Report

### v1

@@ Abstract
Oscar Sainz1 Iker García-Ferrero1 Alon Jacovi2 Jon Ander Campos3Yanai Elazar4,5Eneko Agirre1Yoav Goldberg2,4 Wei-Lin Chen6,7Jenny Chim8Leshem Choshen9,10Luca D’Amico-Wong11 Melissa Dell11Run-Ze Fan12Shahriar Golchin13Yucheng Li14Pengfei Liu12 Bhavish Pahwa15Ameya Prabhu17Suryansh Sharma18Emily Silcock11 Kateryna Solonko15David Stap19Mihai Surdeanu20Yu-Min Tseng21 Vishaal Udandarao16,17,22Zengzhi Wang12Ruijie Xu12Jinglin Yang11 1HiTZ Center - Ixa, University of the Basque Country UPV/EHU  2Bar Ilan University  3Cohere 4Allen Institute for Artificial Intelligence  5University of Washington  6National Taiwan University 7University of Virginia  8Queen Mary University of London  9MIT-IBM Watson AI Lab  10MIT 11Harvard University  12Shanghai Jiao Tong University  13University of Arizona  14University of Surrey 15Microsoft Research  16Tübingen AI Center  17University of Tübingen 18Indian Institute of Technology Kharagpur  19University of Amsterdam  20University of Arizona 21Microsoft  22University of Cambridge Contact: conda-workshop@googlegroups.com

### v2

# Data Contamination Report from the 2024 CONDA Shared Task

@@ Abstract
Oscar Sainz1 Iker García-Ferrero1 Alon Jacovi2 Jon Ander Campos3Yanai Elazar4,5Eneko Agirre1Yoav Goldberg2,4 Wei-Lin Chen6,7Jenny Chim8Leshem Choshen9,10Luca D’Amico-Wong11 Melissa Dell11Run-Ze Fan12Shahriar Golchin13Yucheng Li14Pengfei Liu12 Bhavish Pahwa15Ameya Prabhu16,17Suryansh Sharma18Emily Silcock11 Kateryna Solonko15David Stap19Mihai Surdeanu13Yu-Min Tseng6 Vishaal Udandarao17,21Zengzhi Wang12Ruijie Xu12Jinglin Yang11 1HiTZ Center - Ixa, University of the Basque Country UPV/EHU  2Bar Ilan University  3Cohere 4Allen Institute for Artificial Intelligence  5University of Washington  6National Taiwan University 7University of Virginia  8Queen Mary University of London  9MIT-IBM Watson AI Lab  10MIT 11Harvard University  12Shanghai Jiao Tong University  13University of Arizona  14University of Surrey 15Microsoft Research  16Tübingen AI Center  17University of Tübingen 18Indian Institute of Technology Kharagpur  19University of Amsterdam  20Microsoft  21University of Cambridge Contact: conda-workshop@googlegroups.com